What is the percent composition, to two significant digits, for each element in sodium chloride (NaCl)?
1
Expert's answer
2020-12-18T08:45:31-0500

Firstly we meed to find the molar mass

Molar mass of NaCl=23+35.5=58.5

The Percent composition are as follows;

Na=(23/58.5)×100=39%

Cl=(35.5/58.5)×100=61%